Solve the PDE ∂z/∂x + ∂z/∂y = z2

1
Expert's answer
2021-03-10T09:09:18-0500

"\frac{\partial z}{\partial x}+\frac{\partial z}{\partial y}=z^2\\
\text{The auxillairy equation are, }\\
dx=dy=\frac{dz}{z^2}\\
\text{Integrate the first two fractions. We have that; }\\
x=y+a\\
x-y=a\\
\text{Integrate the last two fractions }\\
y=-\frac{1}{z}+b\\
y+\frac{1}{z}=b.\\
\text{Hence the solution is } f(a,b)=f(x-y,y+\frac{1}{z})=0."